# Bridge multisections of knotted surfaces in $S^{4}$

### v1

@@ 3. Bridge multisections
Only the unknotted 2–sphere admits a one–bridge multisection. In this section, we aim to determine all two–bridge multisected surfaces in 4–space (Thm 3.6). This argument follows the same outline as the work in [IKLCM22, Sec 5.3]. We include a brief commentary for completeness.
Trivial tangles with two bridges are often called rational tangles as they are in correspondence with the extended rational numbers $\mathbb{Q}\cup\{\frac{1}{0}\}$ [KL04]. The following result is a consequence of Conway’s classification of 2–bridge links.
@@
    $T_{1}\cup\overline{T}_{2}$ is an unknot if an only if $i\left(\frac{p_{1}}{q_{1}},\frac{p_{2}}{q_{2}}\right)=|p_{1}q_{2}-p_{2}q_{1}|=1$.
By Lemma 3.4, a 2–bridge multiplane diagram $\mathcal{T}$ can be represented by a finite sequence of extended rational numbers $(\alpha_{1},\dots,\alpha_{n})$ where $i(\alpha_{i},\alpha_{i+1})=1$ or $\alpha_{i}=\alpha_{i+1}$ for each $i=1,\dots,n$. Sequences of this form are loops (with possibly constant edges) in the Farey graph (see [Hat22] for more details).
Lemma 3.5 (Prop 5.10 of [IKLCM22]).
@@
We will proceed by induction in the number of sectors of $\mathcal{T}$. By [MZ17, Sec 4.3], the result holds for $n$–sections with $n\leq 3$. Assume that the result holds some $n\geq 3$ and let $\mathcal{T}$ be a 2–bridge $(n+1)$–section. If $F_{\mathcal{T}}$ is disconnected, then all the tangles of $\mathcal{T}$ must be the same and $F$ is a 2–component unlink of 2–spheres. We assume that $F_{\mathcal{T}}$ is connected and the tuple $\mathcal{T}$ is not constant. If $\mathcal{T}$ has a pair of consecutive tangles with $T_{i}=T_{i+1}$, then the disk $\mathcal{D}_{i}$ between them is isotopic to a product disk. Thus, erasing $T_{i}$ from $\mathcal{T}$ would not affect the underlying surface. By the inductive hypothesis, the shorter tuple represents an unknotted surface and so $F_{\mathcal{T}}$ is unknotted.
Assume that no consecutive tangles of $\mathcal{T}$ are equal. By Lemma 3.5, after a cyclic permutation of $\mathcal{T}$, we can choose a tuple of extended rational numbers representing $\mathcal{T}$ such that the first two numbers are $\alpha_{1}=0/1$ and $\alpha_{2}=1/0$ and $\alpha_{3}$ is ±$\pm 1$ or $0/1$. In each of these cases, $T_{3}\cup\overline{T}_{1}$ is an unlink and the tuple $\mathcal{T}_{0}=(T_{1},T_{2},T_{3})$ is a triplane diagram. By the same reason, removing $T_{2}$ from $\mathcal{T}$ still yields an $n$–plane diagram $\mathcal{T}^{\prime}=(T_{1},T_{3},T_{4},\dots,T_{n+1})$. By the inductive hypothesis, $\mathcal{T}_{0}$ and $\mathcal{T}^{\prime}$ represent unknotted surfaces in 4–space. If $\alpha_{3}$ is ±$\pm 1$, $T_{3}\cup\overline{T}_{1}$ is an unknot and the 3–sphere $B_{3}\cup\overline{B}_{1}$ is a connect summing sphere for $F_{\mathcal{T}}$ with connect summands the surfaces $F_{\mathcal{T}_{0}}$ and $F_{\mathcal{T}^{\prime}}$. Hence, $F_{\mathcal{T}}=F_{\mathcal{T}_{0}}\#F_{\mathcal{T}^{\prime}}$ is also unknotted. If $\alpha_{3}$ is $0/1$, the link $T_{3}\cup\overline{T}_{1}$ is a 2–component unlink and the triplane diagram $\mathcal{T}_{0}=(T_{1},T_{2},T_{3})$ represents an unknotted 2–sphere. This way, the action of inserting $T_{2}$ into $\mathcal{T}^{\prime}$ corresponds to replacing two disks of $F_{\mathcal{T}^{\prime}}$ with one annulus. In words, $F_{\mathcal{T}}$ is obtained by adding a 1–handle to $F_{\mathcal{T}^{\prime}}$. As consecutive tangles of $\mathcal{T}$ are distinct, $\mathcal{T}^{\prime}$ is not a constant sequence of tangles, and so $F_{\mathcal{T}^{\prime}}$ is a connected unknotted surface. On the other hand, Lemma 2.1 states that adding a 1–handle to a connected unknotted surface yields an unknotted surface. Hence $F_{\mathcal{T}}$ is also unknotted. ∎
@@ 3.2. Crossingless multisections
Below, we ask for a strengthening of this proposition. If the answer to Question 3.8 is positive, one may expect a proof using the technology in Section 4.
Question 3.8.
@@ 3.3. New from old
Given two surfaces $F_{1}$ and $F_{2}$, their distant sum $F_{1}\sqcup F_{2}$ is defined by taking the connected sum of their ambient spaces using small 4–ball neighborhoods of points disjoint from $F_{1}$ and $F_{2}$. If one instead chooses neighborhoods of points in $F_{1}$ and $F_{2}$, the resulting surface in 4–space is $F_{1}\#F_{2}$ the connected sum of $F_{1}$ and $F_{2}$. If the surfaces are equipped with bridge $n$–sections $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$, then one can choose the connecting sum points between the ambient 4–spaces to lie in the central surfaces of $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$. This will yield $(b_{1}+b_{2})$–bridge $n$–sections for $F_{1}\sqcup F_{2}$ and $(b_{1}+b_{2}-1)$–bridge $n$–sections for $F_{1}\#F_{2}$. At the level of multiplane diagrams, this translates as taking the distant sum or the pointed-connected sum of the respective tangles. See [MZ17, Sec 2.2] and [ABG+23, Sec 2.3] for more details.
@@ 3.4. Completely decomposable diagrams
Triplane diagrams with at most three bridges have been classified [MZ17, Thm 1.8]. A completely decomposable triplane diagram is one built out of low-bridge diagrams. More precisely, a triplane diagram is completely decomposable if it can be written as the distant sum of connected sums of triplane diagrams with at most three bridges. These triplane diagrams will play a role when defining the merging move in Section 5.2. The following propositions are sufficient conditions for a triplane diagram to be completely decomposable. In [APZ23, Thm 3.5], it was proven that triplane diagrams with $\mathcal{L}^{*}(\mathcal{T})\leq 2$ are also completely decomposable.
@@
Less is known for completely decomposable bridge multisections with at least four sectors. Given $n\geq 4$, a $b$–bridge $n$–section is completely decomposable if it can be written as the distant sum of the connected sum of multisections with bridge number at most two. In [ABG+23, Thm 6.2], the authors used the $\mathcal{L}^{*}$–invariant to give a criterion for completely decomposable multisections.
@@ 3.5. Multisections of twist spun knots
Let $\beta$ be a $(2b-1)$–strand braid and $m$ a fixed integer. We define the link $K(\beta)$ in $b$–bridge position, and the $(2b-1)$–strand tangles $T_{a}(\beta)$, $T_{b}(\beta,m)$, and $T_{c}(\beta)$ as in Figure 6. In [ABG+23, Sec 4.3], the authors explained how $\left(T_{a}(\beta),T_{b}(\beta,m),T_{c}(\beta),T_{b}(\beta,0)\right)$ is a multiplane diagram for a $(2b-1,b)$–bridge 4–section of the $m$–twist spun of $K(\beta)$. In fact, $\left(T_{a}(\beta),T_{b}(\beta,m),T_{c}(\beta)\right)$ is a bisection for the ribbon disk obtained by $m$–twisting the half spin of $K(\beta)$. The following is a consequence of the observation that the tuple $\left(T_{c}(\beta),T_{b}(\beta,m),T_{a}(\beta)\right)$ represents the same ribbon disk in the opposite orientation.
@@
is a $(2b-1;b)$–multiplane diagram for the $m$–twist spin of $K$. We denote such tuple as $\mathcal{T}\left(\beta;(m_{1},\dots,m_{2l})\right)$.
@@
contains infinitely many non-diffeomorphic 8–sections of the same $k$–twist spun knot. One can see this by comparing the cross-sections of such multiplane diagrams. The following result exploits this idea to build distinct multiplane diagrams of the same surface. Notice that Theorem 3.12 is stated for multisections with at least four sectors. The trisection version ($n=3$) of this result is still open.
@@
Suppose that $F\subset S^{4}$ admits a $b$–bridge $n$–section for some $b\geq 2$, $n\geq 4$. Then for any $b^{\prime}\geq b+2$, $F$ admits infinitely many non-diffeomorphic $b^{\prime}$–bridge $n$–sections.
@@
We will find bridge multisections $\{\mathcal{T}^{p}_{m,n}:p\geq 3,m\geq 1,n\geq 4\}$ satisfying the following properties: (1) each $\mathcal{T}^{p}_{m,n}$ is a $p$–bridge $n$–section for an unknotted 2–sphere and, (2) the cross-sections of $\mathcal{T}^{p}_{m,n}$ are the split union of unlinks and at most one $(3,3m)$–torus link. This way, if $\mathcal{T}$ is a $b$–bridge $n$–section for $F$ for some $n\geq 4$, then $\mathcal{F}^{p}_{n}=\{\mathcal{T}\#\mathcal{T}^{p}_{m,n}:m\geq 1\}$ will be a collection of $(b+p-1)$–bridge $n$–sections representing $F$.
We explain how to build $\mathcal{T}^{3}_{m,n}$. Let $\sigma_{1},\sigma_{2}\in B_{3}$ be the standard 3-braid generators and let $\vec{V}_{m}$ be the vector $(m,0,0,\dots,0)\in\mathbb{Z}^{2l}$. As $K(\sigma_{1})$ is a $2$–bridge unknot, Proposition 3.11 states that $\mathcal{T}^{3}_{m,4l}=\mathcal{T}(\sigma_{1},\vec{V}_{m})$ is a $3$–bridge $4l$–section of the unknotted 2–sphere. By construction, all the cross-sections of $\mathcal{T}^{3}_{m,4l}$ are of the form $T_{a}(\sigma_{1})\cup\overline{T_{b}(\sigma_{1},m)}$, $T_{a}(\sigma_{1})\cup\overline{T_{b}(\sigma_{1},0)}$, $T_{a}(\sigma_{1})\cup\overline{T_{c}(\sigma_{1})}$, $T_{c}(\sigma_{1})\cup\overline{T_{b}(\sigma_{1},m)}$, $T_{c}(\sigma_{1})\cup\overline{T_{b}(\sigma_{1},0)}$, and $T_{b}(\sigma_{1},m)\cup\overline{T_{b}(\sigma_{1},0)}$. Observe that the first five links are unlinks, and the last is the $(3,3m)$–torus link. To build $\mathcal{T}^{3}_{m,n}$ for $n$ not a multiple of four, we can add trivial sectors by duplicating the desired amount of tangles in our multiplane diagram $\mathcal{T}^{3}_{m,4l}$.
@@
We now explain why the collection $\mathcal{F}^{p}_{n}$ contains infinitely many pairwise non-diffeomorphic $n$–sections. Notice first that $(3,3m)$–torus links yield non-isotopic links for distinct choices of integers $m\geq 1$ and that the cross-sections of $\mathcal{T}\#\mathcal{T}^{p}_{m,n}$ are connected sums of cross-sections of $\mathcal{T}$ and $\mathcal{T}^{p}_{m,n}$. These observations, together with the uniqueness of prime decompositions for links [Has58], imply the existence of an infinite set $\mathcal{G}^{p}_{n}\subset\mathcal{F}^{p}_{n}$ with the following property: for each pair of distinct $x,y\in\mathcal{G}^{p}_{n}$, at least one cross-section of $x$ is not isotopic to any cross-section of $y$. Hence, by Lemma 3.3, $\mathcal{G}^{p}_{n}$ is an infinite set of pairwise non-diffeomorphic $n$–sections for $F$. ∎

### v2

@@ 3. Bridge multisections
Only the unknotted 2–sphere admits a one–bridge multisection. In this section, we aim to determine all two–bridge multisected surfaces in $S^{4}$ (Thm 3.7). This argument follows the same outline as the work in [IKL+22, Sec 5.3]. We include a brief commentary for completeness.
Trivial tangles with two bridges are often called rational tangles as they are in correspondence with the extended rational numbers $\mathbb{Q}\cup\{\frac{1}{0}\}$ [KL04]. For a pair of extended rational numbers, let $i\left(\frac{p_{1}}{q_{1}},\frac{p_{2}}{q_{2}}\right)$ be equal to $|p_{1}q_{2}-p_{2}q_{1}|$. The following result is a consequence of Conway’s classification of 2–bridge links.
@@
    $T_{1}\cup\overline{T}_{2}$ is an unknot if and only if $i\left(\frac{p_{1}}{q_{1}},\frac{p_{2}}{q_{2}}\right)=1$.
By Lemma 3.5, a 2–bridge multiplane diagram $\mathcal{T}$ can be represented by a finite sequence of extended rational numbers $(\alpha_{1},\dots,\alpha_{n})$ where $i(\alpha_{i},\alpha_{i+1})=1$ or $\alpha_{i}=\alpha_{i+1}$ for each $i=1,\dots,n$. Sequences of this form are loops (with possibly constant edges) in the Farey graph (see [HAT02] for more details).
Lemma 3.6 (Prop 5.10 of [IKL+22]).
@@
We will proceed by induction in the number of sectors of $\mathcal{T}$. By [MZ17, Sec 4.3], the result holds for nn–sections with $n=3$. Assume that the result holds for some $n\geq 3$ and let $\mathcal{T}$ be a 2–bridge (n+1)(n+1)–section. If $F_{\mathcal{T}}$ is disconnected, then all the tangles of $\mathcal{T}$ must be the same and $F$ is a 2–component unlink of 2–spheres. We assume that $F_{\mathcal{T}}$ is connected and the $\mathcal{T}$ is not a constant sequence of tangles. If $\mathcal{T}$ has a pair of consecutive tangles with $T_{i}=T_{i+1}$, then the disks $\mathcal{D}_{i}$ between them are isotopic to product disks. Thus, erasing $T_{i}$ from $\mathcal{T}$ would not affect the underlying surface. By the inductive hypothesis, the shorter tuple represents an unknotted surface and so $F_{\mathcal{T}}$ is unknotted.
Assume that no consecutive tangles of $\mathcal{T}$ are equal. By Lemma 3.6, after a cyclic permutation of $\mathcal{T}$, we can choose a tuple of extended rational numbers representing $\mathcal{T}$ such that the first two numbers are $\alpha_{1}=0/1$ and $\alpha_{2}=1/0$ and $\alpha_{3}$ is $\pm 1$ or $0/1$. In these cases, Lemma 3.5 implies that $T_{3}\cup\overline{T}_{1}$ is an unlink and the tuple $\mathcal{T}_{0}=(T_{1},T_{2},T_{3})$ is a triplane diagram. By the same reason, removing $T_{2}$ from $\mathcal{T}$ still yields an nn–plane diagram $\mathcal{T}^{\prime}=(T_{1},T_{3},T_{4},\dots,T_{n+1})$. By the inductive hypothesis, $\mathcal{T}_{0}$ and $\mathcal{T}^{\prime}$ represent unknotted surfaces in $S^{4}$. If $\alpha_{3}$ is $\pm 1$, $T_{3}\cup\overline{T}_{1}$ is an unknot and the 3–sphere $B_{3}\cup\overline{B}_{1}$ is a connect summing sphere for $F_{\mathcal{T}}$ with connect summands the surfaces $F_{\mathcal{T}_{0}}$ and $F_{\mathcal{T}^{\prime}}$. Hence, $F_{\mathcal{T}}=F_{\mathcal{T}_{0}}\#F_{\mathcal{T}^{\prime}}$ is also unknotted. If $\alpha_{3}$ is $0/1$, the link $T_{3}\cup\overline{T}_{1}$ is a 2–component unlink and the triplane diagram $\mathcal{T}_{0}=(T_{1},T_{2},T_{3})$ represents an unknotted 2–sphere. This way, the action of inserting $T_{2}$ into $\mathcal{T}^{\prime}$ corresponds to replacing two disks of $F_{\mathcal{T}^{\prime}}$ with one annulus. In words, $F_{\mathcal{T}}$ is obtained by adding a 1–handle to $F_{\mathcal{T}^{\prime}}$. As consecutive tangles of $\mathcal{T}$ are distinct, $\mathcal{T}^{\prime}$ is not a constant sequence of tangles, and so $F_{\mathcal{T}^{\prime}}$ is a connected unknotted surface. On the other hand, Lemma 2.1 states that adding a 1–handle to a connected unknotted surface yields an unknotted surface. Hence $F_{\mathcal{T}}$ is also unknotted. ∎
@@ 3.2. Crossingless multisections
Proposition 3.8 was proven for triplane diagrams in [MZ17, Prop 4.4]. That said, the same proof works for arbitrary nn–plane diagrams. Below, we ask for a strengthening of this proposition. If the answer to Question 3.9 is positive, one may expect a proof using the technology in Section 4.
Question 3.9.
@@ 3.3. New from old
Given two surfaces $F_{1}$ and $F_{2}$, their distant sum $F_{1}\sqcup F_{2}$ is defined by taking the connected sum of their ambient spaces using small 4–ball neighborhoods of points disjoint from $F_{1}$ and $F_{2}$. If one instead chooses neighborhoods of points in $F_{1}$ and $F_{2}$, the resulting surface in $S^{4}$ is $F_{1}\#F_{2}$ the connected sum of $F_{1}$ and $F_{2}$. If the surfaces are equipped with bridge nn–sections $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$, then one can choose the connecting sum points between the ambient $S^{4}$s to lie in the central surfaces of $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$. This will yield $(b_{1}+b_{2})$–bridge nn–sections for $F_{1}\sqcup F_{2}$ and $(b_{1}+b_{2}-1)$–bridge nn–sections for $F_{1}\#F_{2}$. At the level of multiplane diagrams, this translates as taking the distant sum or the pointed-connected sum of the respective tangles. See [MZ17, Sec 2.2] and [ABG+23a, Sec 2.3] for more details.
@@ 3.4. Completely decomposable diagrams
Triplane diagrams with at most three bridges have been classified [MZ17, Thm 1.8]. A completely decomposable bridge trisection is one built out of low-bridge trisections. More precisely, a bridge trisection is completely decomposable if it can be written as the distant sum of connected sums of bridge trisections with at most three bridges. These bridge trisections will play a role when defining the merging move in Section 5.2. The following propositions are sufficient conditions for a bridge trisection to be completely decomposable. In [APZ23b, Thm 3.5], it was proven that bridge trisections with $\mathcal{L}^{*}(\mathcal{T})\leq 2$ are also completely decomposable.
@@
Less is known about completely decomposable bridge multisections with at least four sectors. Given $n\geq 4$, a bb–bridge nn–section is completely decomposable if it can be written as the distant sum of the connected sum of multisections with bridge number at most two. In [ABG+23a, Thm 6.2], the authors used the $\mathcal{L}^{*}$–invariant to give a criterion for completely decomposable multisections.
For the invested reader, the collection $\{\mathcal{T}^{p}_{m,n}:p\geq 3,m\geq 1,n\geq 4\}$ in the proof of Theorem 3.13 contains non-completely decomposable nn–sections of the unknotted 2–sphere for all $n\geq 4$. This claim is true as the cross-sections of completely decomposable multisections are distant sums of unlinks and connected sums of 2–bridge links.
@@ 3.5. Multisections of twist spun knots
In this subsection, we will use the concept of bridge bisection of a surface with boundary properly embedded in $B^{4}$, which we define now. Suppose $T_{a}$, $T_{b}$, and $T_{c}$ are bb–bridge trivial tangles with the same endpoints such that $T_{a}\cup\overline{T}_{b}$ and $T_{b}\cup\overline{T}_{c}$ are unlinks but $T_{a}\cup\overline{T}_{c}$ may be knotted. Let $(B^{4},\mathcal{D}_{ab})$ and $(B^{4},\mathcal{D}_{bc})$ be trivial disks bounded by $T_{a}\cup\overline{T}_{b}$ and $T_{b}\cup\overline{T}_{c}$. A bridge bisection for a surface $F\subset B^{4}$ is a decomposition $(B^{4},\mathcal{D}_{ab})\cup_{(B^{3},T_{b})}(B^{4},\mathcal{D}_{bc})$. In this case, we say that $(T_{a},T_{b},T_{c})$ is a bisection diagram for $F$. Examples of bridge bisections for surfaces with boundary can be built by removing one 4-dimensional sector from a bridge trisection of a closed surface. The proof of Theorem 1.3 in [MZ17] can be used to show that every ribbon surface in $B^{4}$ admits a bridge bisection; we leave it as an exercise for the reader to verify this claim.
Let $\beta$ be a (2b1)(2b-1)–strand braid and $m$ a fixed integer. We define the link $K(\beta)$ in bb–bridge position, and the (2b1)(2b-1)–strand tangles $T_{a}(\beta)$, $T_{b}(\beta,m)$, and $T_{c}(\beta)$ as in Figure 6. In [ABG+23a, Sec 4.3], the authors explained how $\left(T_{a}(\beta),T_{b}(\beta,m),T_{c}(\beta),T_{b}(\beta,0)\right)$ is a multiplane diagram for a (2b1,b)(2b-1,b)–bridge 4–section of the mm–twist spun of $K(\beta)$. In fact, $\left(T_{a}(\beta),T_{b}(\beta,m),T_{c}(\beta)\right)$ is a bridge bisection diagram for the ribbon disk obtained by mm–twisting the half spin of $K(\beta)$. The following is a consequence of the observation that the tuple $\left(T_{c}(\beta),T_{b}(\beta,m),T_{a}(\beta)\right)$ represents the same ribbon disk in the opposite orientation.
@@
is a (2b1;b)(2b-1;b)–multiplane diagram for the mm–twist spin of $K$. We denote such tuple as $\mathcal{T}\left(\beta;(m_{1},\dots,m_{2l})\right)$.
@@
contains infinitely many non-diffeomorphic 8–sections of the same kk–twist spun knot. One can see this by comparing the cross-sections of such multiplane diagrams. The following result exploits this idea to build distinct multiplane diagrams of the same surface. Notice that Theorem 3.13 is stated for multisections with at least four sectors. The trisection version ($n=3$) of this result is still open.
@@
Suppose that $F\subset S^{4}$ admits a bb–bridge nn–section for some $b\geq 2$, $n\geq 4$. Then for any $b^{\prime}\geq b+2$, $F$ admits infinitely many non-diffeomorphic $b^{\prime}$–bridge nn–sections.
@@
We will find bridge multisections $\{\mathcal{T}^{p}_{m,n}:p\geq 3,m\geq 1,n\geq 4\}$ satisfying the following properties: (1) each $\mathcal{T}^{p}_{m,n}$ is a pp–bridge nn–section for an unknotted 2–sphere and, (2) the cross-sections of $\mathcal{T}^{p}_{m,n}$ are the split union of unlinks and at most one (3,3m)(3,3m)–torus link. This way, if $\mathcal{T}$ is a bb–bridge nn–section for $F$ for some $n\geq 4$, then $\mathcal{F}^{p}_{n}=\{\mathcal{T}\#\mathcal{T}^{p}_{m,n}:m\geq 1\}$ will be a collection of (b+p1)(b+p-1)–bridge nn–sections representing $F$.
We explain how to build $\mathcal{T}^{3}_{m,n}$. Let $\sigma_{1},\sigma_{2}\in B_{3}$ be the standard 3–braid generators and let $\vec{V}_{m}$ be the vector $(m,0,0,\dots,0)\in\mathbb{Z}^{2l}$. As $K(\sigma_{1})$ is a 22–bridge unknot, Proposition 3.12 states that $\mathcal{T}^{3}_{m,4l}=\mathcal{T}(\sigma_{1},\vec{V}_{m})$ is a 33–bridge 4l4l–section of the unknotted 2–sphere. By construction, all the cross-sections of $\mathcal{T}^{3}_{m,4l}$ are of the form $T_{a}(\sigma_{1})\cup\overline{T_{b}(\sigma_{1},m)}$, $T_{a}(\sigma_{1})\cup\overline{T_{b}(\sigma_{1},0)}$, $T_{a}(\sigma_{1})\cup\overline{T_{c}(\sigma_{1})}$, $T_{c}(\sigma_{1})\cup\overline{T_{b}(\sigma_{1},m)}$, $T_{c}(\sigma_{1})\cup\overline{T_{b}(\sigma_{1},0)}$, and $T_{b}(\sigma_{1},m)\cup\overline{T_{b}(\sigma_{1},0)}$. Observe that the first five links are unlinks, and the last is the (3,3m)(3,3m)–torus link. To build $\mathcal{T}^{3}_{m,n}$ for $n$ not a multiple of four, we can add trivial sectors by duplicating the desired amount of tangles in our multiplane diagram $\mathcal{T}^{3}_{m,4l}$.
@@
We now explain why the collection $\mathcal{F}^{p}_{n}$ contains infinitely many pairwise non-diffeomorphic nn–sections. Notice first that (3,3m)(3,3m)–torus links yield non-isotopic links for distinct choices of integers $m\geq 1$ and that the cross-sections of $\mathcal{T}\#\mathcal{T}^{p}_{m,n}$ are connected sums of cross-sections of $\mathcal{T}$ and $\mathcal{T}^{p}_{m,n}$. These observations, together with the uniqueness of prime decompositions for links [HAS58], imply the existence of an infinite set $\mathcal{G}^{p}_{n}\subset\mathcal{F}^{p}_{n}$ with the following property: for each pair of distinct $x,y\in\mathcal{G}^{p}_{n}$, at least one cross-section of $x$ is not isotopic to any cross-section of $y$. Hence, by Lemma 3.4, $\mathcal{G}^{p}_{n}$ is an infinite set of pairwise non-diffeomorphic nn–sections for $F$. ∎